Q9) The Almarai Co. just paid a dividend of $1.95 per share on its stock. The dividends are to grow at a constant rate of 6 percent per year indefinitely. If investors require an 1I percent return on Almarai Co. stock, what is the current price? What will the priee be in three years? In 15 years? 5/6
Answer #1
A B C D E F G H I J K
2
3 As per dividend growth model, Price of share is the present value of all future dividends discounted at cost of equity.
4 Current Price of Share (P) =Div0*(1+g)/(r-g)
5 Where Div0 is the dividend paid this year, r is cost of equity and g is perpetual growth rate of dividend.
6
7 Given the following data:
8 Dividend (Div0) $1.95
9 Cost of equity (r) 11%
10 Growth rate (g) 6.0%
11
12 From Dividend growth model,
13 Share Price, P =Div0*(1+g)/(r-g)
14 $41.34 =D8*(1+D10)/(D9-D10)
15
16 Hence Current Share Price is $41.34
17
18 Calculation of Share Price at Year 3:
19
20 Given the following data:
21 Dividend (Div3) $2.32 =D8*((1+D10)^3)
22 Cost of equity (r) 11%
23 Growth rate (g) 6.0%
24
25 From Dividend growth model,
26 Share Price at year 3, P3 =Div3*(1+g)/(r-g)
27 $49.24 =D21*(1+D23)/(D22-D23)
28
29 Hence Share Price at Year 3 is $49.24
30
31 Calculation of Share Price at Year 15:
32
33 Given the following data:
34 Dividend (Div15) $4.67 =D8*((1+D23)^15)
35 Cost of equity (r) 11%
36 Growth rate (g) 6.0%
37
38 From Dividend growth model,
39 Share Price at year 15, P15 =Div15*(1+g)/(r-g)
40 $99.07 =D34*(1+D36)/(D35-D36)
41
42 Hence Share Price at Year 15 is $99.07
43